OpenStudy (anonymous):

16x-15-9x=13 solve the equation someone please help and sho how pleaseeee

OpenStudy (anonymous):

combine like terms first, then separate variable terms from numerical terms.

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

7x-15=13 7x=28 x=28/7 x=4

OpenStudy (anonymous):

oops nvmd

OpenStudy (anonymous):

howd u get 28?

OpenStudy (anonymous):

adding 15 to both sides. 7x-15(+15)=13(+15) 7x=28

OpenStudy (anonymous):

my teacher said to check it how do i check if it is right?

OpenStudy (anonymous):

answer is x=4. So you replace the value of x with 4 in the original equation. i.e. 16x-15-9x=13 16(4)-15-9(4)=13 64-15-36=13 13=13 Therefore its correct that the solution for 16x-15-9x=13, x=4

OpenStudy (anonymous):

thank hyou